1. write an equation of the line containing (6, 5) and having slope 3/4.
a. express the equation in f(x) = mx + b form
b. create a table of values with at least 5 ordered pairs.
c. give the domain, range, x- and y-intercepts and trend.
d. sketch the graph of the line.
e. give the equation of its inverse g(x) = mx + b form
2. 2. Write an equation of the line containing (1, -3) and (-4, -2).
a. Express the equation in f(x) = mx + b form.
b. Create a table of values with at least 5 ordered pairs.
c. Give the domain, range, slope, x-and y-intercepts and trend.
d. Sketch the graph of the line.
e. Give the equation of its inverse g(x) = mx + b form.
3. Write an equation of the line whose x-intercept is 5 and y-intercept is -3.
a. Express the equation in f(x) = mx + b form.
b. Create a table of values with at least 5 ordered pairs.
c. Give the domain, range, slope, and trend.
d. Sketch the graph of the line.
e. Give the equation of its inverse g(x) = mx + b form.
4. Write an equation of the line which is parallel to 5x – 3y = 15 that passes through (2, 3).
a. Express the equation in f(x) = mx + b form.
b. Create a table of values with at least 5 ordered pairs.
c. Give the domain, range, slope, x-and y-intercepts and trend.
d. Sketch the graph of the line.
e. Give the equation of its inverse g(x) = mx + b form.
5. Write an equation of the line which is perpendicular to 5x – 3y = 15 that passes through (2, 3).
a. Express the equation in f(x) = mx + b form.
b. Create a table of values with at least 5 ordered pairs.
c. Give the domain, range, slope, x-and y-intercepts and trend.
d. Sketch the graph of the line.
e. Give the equation of its inverse g(x) = mx + b form.

Answer by oscargut(2103)

Hi, i am doing 1) here
1. write an equation of the line containing (6, 5) and having slope 3/4.
a. express the equation in f(x) = mx + b form
b. create a table of values with at least 5 ordered pairs.
c. give the domain, range, x- and y-intercepts and trend.
d. sketch the graph of the line.
e. give the equation of its inverse g(x) = mx + b form

a) y = (3/4)x+1/2
b)
(0,1/2), (1,5/4), (2,2), (3,11/4) , (4,7/2)
c) domain = range = real numbers , trend is increasing
d) graph the line using (0,1/2) and (2,2)
e) g(x) = (4/3)x-2/3

If you need more help today, you can contact me now at:
mthman@gmail.com
Thanks